OpenStudy (anonymous):

How can this multiplication problem be written using only improper fractions? five times two and one-eighth equals what value? five over one times sixteen over eight equals what value? five over one times eight over sixteen equals what value? five over one times seventeen over eight equals what value? five over one times eighteen over seven equals what value?

OpenStudy (anonymous):

5x2 and 1/8 is first \[5x2\frac{ 1 }{ 8 }\]

OpenStudy (texaschic101):

5 * 2 1/8 = 5/1 * 17/8

OpenStudy (anonymous):

thank you @texaschic101

OpenStudy (texaschic101):

do you understand, because I can explain it better if you don't ?

OpenStudy (anonymous):

Lol sorry my response wasn't right

OpenStudy (anonymous):

i do understand

OpenStudy (texaschic101):

any whole number can be turned into an improper fraction by putting it over 1. Example : 3 = 3/1 or 7 = 7/1. If you have a mixed number and you want to turn it into an improper fraction , like 2 1/8, you take the whole number, multiply it by the denominator and add it to the numerator, then put it over the original denominator. 2 1/8 = (2 * 8 + 1)/8 = 17/8 or 3 1/6 = (3 * 6 + 1)/6 = 19/6 hope that helped